# HaSPeR: An Image Repository for Hand Shadow Puppet Recognition

### v1

@@ 5.3 Error Analysis
The confusion matrix for the InceptionV3 model on our dataset, presented in Figure 11, reveals that the ‘Panther’ class exhibits the highest count of misclassifications. One obvious reason for this is the somewhat significant inter-class similarity among the ‘Dog’, ‘Elephant’, and ‘Panther’ classes. Most of the misclassified samples are from visually similar classes. We can posit that navigating the intricacies of visually similar classes poses a significant challenge in this image classification task, as evident from the other pale-red entries of the confusion matrix in Figure 11. Even to the keen human eye, distinguishing between these classes may be perplexing, as they share common visual features, shapes, or color patterns that result in a high degree of resemblance. We examine various aspects, such as the distinctive features or characteristics that might have led to confusion and the degree of similarity between the misclassified classes. Figure 10(a) and 10(b) show the confusion between a ‘Crab’ sample and a ‘Rabbit’ sample which look visually quite similar. The same holds for the sole ‘Moose’ sample that is misclassified as a ’Bird’ sample by the InceptionV3 model, as depicted in Figure 10(c) and 10(d). We observe that misclassifications of this type occur when images belonging to different, but visually akin categories, are erroneously assigned to the wrong class. The green entries along the diagonal of the confusion matrix in Figure 11 indicate the reasonably good classwise prediction performance of the InceptionV3 model, even though there exists a slight class imbalance in HaSPeR.
@@ 6 Conclusion and Future Work
In this paper, we explore an intriguing subject matter, hand shadow puppetry, in the realm of computer vision. We introduce HaSPeR, a novel dataset with a sizable collection of 8,340 hand shadow puppet images distributed across 11 classes, taken from both expert and amateur puppetry performance clips. We fine-tune 31 pretrained image classification models on HaSPeR to establish a benchmark for the dataset. To explain and analyze the performance of the most erudite model, InceptionV3, in comparison with other baseline models, we visually manifest its feature space using the $t$-SNE dimensionality reduction technique. We also perform thorough qualitative and error analyses for the InceptionV3 model. We envisage the possibility of developing applications for imparting the art of shadowgraphy, via mobile and embedded devices. We claim that this work is novel and significant since it is the first publicly available dataset and study on image classification benchmarking that focuses only on ombromanie. There are many avenues in our work that warrant further investigation. We hope to reconcile those desiderata by enriching our dataset with numerous permutations of arm positions and finger movements, preferably by employing more skilled individuals with varying palm and wrist structures, thereby creating more diverse silhouettes. We stipulate that the dataset, in a more supplemented state, will be better suited to fine-tune models with a very large number of parameters. We also plan to experiment with a gesture detection technology such as MediaPipe [^11] MediaPipe — https://developers.google.com/mediapipe or Microsoft Kinect [^12] Kinect for Windows — https://learn.microsoft.com/en-us/windows/apps/design/devices/kinect-for-windows for leveraging depth coordinates of hand landmarks [58], and assess their efficacy in classifying hand shadow puppets.
@@ 7 Acknowledgments
We convey our heartfelt gratitude, in advance, to the anonymous reviewers for their constructive criticisms and insightful feedback which will surely be conducive to the improvement of the research work outlined in this paper. We also appreciate the Systems and Software Lab (SSL) of the Islamic University of Technology (IUT) for the generous provision of computing resources during the course of this project. Additionally, we wish to acknowledge Shahriar Ivan, Department of Computer Science and Engineering, Islamic University of Technology, for his invaluable assistance in proofreading and offering a preliminary review of this manuscript. Syed Rifat Raiyan, in particular, wants to thank his parents, Syed Sirajul Islam and Kazi Shahana Begum, for everything.

### v3

@@ 5.3 Error Analysis
The confusion matrix for the ResNet34 model on our dataset, presented in Figure 12, reveals that the ‘Crab’ class exhibits the highest count of misclassifications. One obvious reason for this is the somewhat significant inter-class similarity among the ‘Bird’, ‘Moose’, ‘Rabbit’, and ‘Crab’ classes. Most of the misclassified samples are from visually similar classes. We can posit that navigating the intricacies of visually similar classes poses a significant challenge in this image classification task, as evident from the other pale-red entries of the confusion matrix in Figure 12. Even to the keen human eye, distinguishing between these classes may be perplexing, as they share common visual features, shapes, or color patterns that result in a high degree of resemblance. We examine various aspects, such as the distinctive features or characteristics that might have led to confusion and the degree of similarity between the misclassified classes. Figure 11(a) and 11(f) show the confusion between a ‘Crab’ sample and a ‘Rabbit’ sample which look visually quite similar. The same holds for one of the six ‘Moose’ samples that are misclassified as ‘Bird’ samples by the ResNet34 model, as depicted in Figure 11(b) and 11(g). We observe that misclassifications of this type occur when images belonging to different, but visually akin categories, are erroneously assigned to the wrong class. Another reason for misclassifications is the ambiguity of shape present in mid-action frames. For instance, the ‘Bird’ sample in Figure 11(h) is a transition frame between two successive wing flaps. However, due to the presence of this ambiguous sample in the training set of the ‘Bird’ class, the ResNet34 ends up misclassifying the ‘Panther’ sample in Figure 11(c) as a ‘Bird’ sample. The misclassification portrayed by the pair of Figure 11(d) and 11(i) is due to the combination of poor lighting and ineptitude of the amateur child puppeteer in creating ‘Llama’ shadows. The model confuses the ear protrusions of the ‘Llama’ sample to be the tentacular eyes of a ‘Snail’ sample. For Figure 11(e), we can pontificate the misclassification reasons to be the overlapping of the fingers and the distorted angle at which the sample was captured. The ‘Snail’ sample thereby gets wrongly classified as a ‘Bird’ sample due to the existence of the analogous sample portrayed in Figure 11(j). The green entries along the diagonal of the confusion matrix in Figure 12 indicate the reasonably good classwise prediction performance of the ResNet34 model, which is, to some extent, due to the perfectly balanced sample distribution in HaSPeR.
5.4 Avenues of Improvement
To reduce the number of misclassifications, models need to be imbued with the ability to learn certain nuanced features. In light of the contemporary image classification literature, we can opt to use Kolmogorov–Arnold Networks (KAN) [58] instead of a simple Multilayer Perceptron (MLP) as the classifier block. The use of Convolutional KANs [59] has yielded good results in many image classification benchmarks. The task of image classification on HaSPeR can be dubbed as a silhouette classification task, which is why we can leverage topological features of the shadow contours to achieve better results [60]. The silhouette polygonization algorithm (PoG), as proposed by Göçmen and Akata [61], may aid in achieving better classification accuracy. Other possible avenues may involve the use of ensemble methods coupled with voting schemes, or resorting to data augmentation with synthetically generated samples, but we defer the exploration of these hypotheses for future research endeavors.
Figure 12: Confusion Matrix of vanilla ResNet34.
@@ 6 Conclusion and Future Work
In this paper, we explore an intriguing subject matter, hand shadow puppetry, in the realm of computer vision. We introduce HaSPeR, a novel dataset with a sizable collection of 15,000 hand shadow puppet images distributed across 15 classes, taken from both expert and amateur puppetry performance clips. We fine-tune 31 pretrained image classification models on HaSPeR to establish a benchmark for the dataset. To explain and analyze the performance of the most erudite model, ResNet34, in comparison with other baseline models, we visually manifest its feature space using the $t$-SNE dimensionality reduction technique. We also perform thorough qualitative and error analyses for the ResNet34 model. We envisage the possibility of developing applications for imparting the art of shadowgraphy, via mobile and embedded devices. We claim that this work is novel and significant since it is the first publicly available dataset and study on image classification benchmarking that focuses only on ombromanie. There are many avenues in our work that warrant further investigation. We hope to reconcile those desiderata by enriching our dataset with numerous permutations of arm positions and finger movements, preferably by employing more skilled individuals with varying palm and wrist structures, thereby creating more diverse silhouettes. We also plan to experiment with a gesture detection technology such as MediaPipe [^11] MediaPipe — https://developers.google.com/mediapipe or Microsoft Kinect [^12] Kinect for Windows — https://learn.microsoft.com/en-us/windows/apps/design/devices/kinect-for-windows for leveraging depth coordinates of hand landmarks [62], and assess their efficacy in classifying hand shadow puppets.
@@ 7 Acknowledgments
We convey our heartfelt gratitude, in advance, to the anonymous reviewers for their constructive criticisms and insightful feedback which will surely be conducive to the improvement of the research work outlined in this paper. We also appreciate the Systems and Software Lab (SSL) of the Islamic University of Technology (IUT) for the generous provision of computing resources during the course of this project. Additionally, we wish to acknowledge Shahriar Ivan, Department of Computer Science and Engineering, Islamic University of Technology, for his invaluable assistance in proofreading and offering a preliminary review of this manuscript. We further thank Mohammad Ishrak Abedin and Reaz Hassan Joader, from the same department, for lending their assistance, which was instrumental in refining this paper’s illustrations and diagrams. Syed Rifat Raiyan, in particular, wants to thank his parents, Syed Sirajul Islam and Kazi Shahana Begum, for everything.